OpenStudy (anonymous):

Complete the pattern using powers of 5 (i will write the equation hold on)

OpenStudy (anonymous):

\[ \frac{ 1 }{ 5^2} = (blank)\]

OpenStudy (anonymous):

\[\frac{ 1 }{ 5^1 } = (blank)\]

OpenStudy (anonymous):

\[\frac{ 1 }{ 5^0} = (blank)\]

OpenStudy (anonymous):

\[\frac{ 1 }{ 5^-1} = (blank)\]

OpenStudy (anonymous):

\[\frac{ 1 }{ 5^-2 } = (blank)\]

OpenStudy (anonymous):

thats the pattern ^^

OpenStudy (anonymous):

How many more do you need?

OpenStudy (anonymous):

no you have to fill in in the blanks @Basher